# Property Enhanced Instruction Tuning for Multi-task Molecule Generation with Large Language Models

### v2

@@ Abstract
Xuan Lin1, Long Chen1, Yile Wang2††thanks: Corresponding author., Xiangxiang Zeng3, Philip S. Yu4 1School of Computer Science, Xiangtan University 2College of Computer Science and Software Engineering, Shenzhen University 3College of Information Science and Engineering, Hunan University 4Department of Computer Science, University of Illinois jack_lin@xtu.edu.cn wangyile@szu.edu.cn psyu@cs.uic.edu
@@ 1 Introduction
Large language models (LLMs) such as GPT-4 OpenAI (2023), PaLM Chowdhery et al. (2023) and LLaMa Touvron et al. (2023); Dubey et al. (2024) have revolutionized the landscape of artificial intelligence and natural language processing (NLP), allowing machines to understand and generate human language with remarkable fluency and coherence. Based on encoded world knowledge Petroni et al. (2019) and powerful instruct-following Zhang et al. (2023) capabilities of LLMs, recent work has successfully used LLM for molecular-related tasks, achieving promising results Fang et al. (2023); Zhang et al. (2024a).
Despite the success, LLMs still have limitations in tasks involving the generation of molecules with restricted properties, therefore limiting its potential applications such as drug discovery Zhavoronkov (2018); Elton et al. (2019). The challenges for tackling such tasks mainly lie in three aspects: (1) Existing studies have shown limitations of LLMs in understanding molecular representations Grisoni (2023), which makes it more challenging for handling such tasks with precise properties; (2) While there is some known SMILES-property pairing data, it often remains limited to predicting a single property and lacks datasets encompassing a wide range of properties Wu et al. (2018). Moreover, most of these datasets do not include precisely described textual data, making it challenging to identify accurate tri-modal data pairs Krenn et al. (2020); (3) To our knowledge, there are no suitable datasets or evaluation methods for multi-constraint molecule generation using LLMs, which poses challenges in standardizing and assessing such molecule generation tasks with these models Jin et al. (2018); Elton et al. (2019).
@@
Our contributions can be summarized as follows:
1) We propose PEIT, a novel framework that enables LLMs to align the textual descriptions, SMILES sequences, and biochemical properties through multi-modal representation alignment, thereby facilitating multi-task molecule generation.
2) We propose a novel molecular design task called multi-constraint molecule generation, simulating the real drug discovery process by setting multiple property constraints to guide LLMs in generating molecules that meet specific requirements. The property values of output molecules can be verified using RDKit Landrum et al. (2013).
3) PEIT achieves promising results in various benchmarks. It surpasses baselines by 2.3% on BLEU-2 in molecule captioning, showing a advantage of 21.76 Levenshtein over baselines in text-based molecule generation, giving best results in five-property constraint molecule generation.
@@ 2 Related Work
Instruction tuning. Specialized datasets construction seems the effective way to enable LLMs to better perform the molecular-related tasks. For instance, Mol-Instructions Fang et al. (2023) provides a large-scale biomolecular instruction dataset designed for LLMs, which contains a variety of instruction data ranging from small molecules, proteins, and biomolecular texts. MolReGPT Li et al. (2023) generates a specialized instruction dataset for chemical reaction prediction and molecular synthesis tasks by integrating molecular structure information with relevant chemical reaction descriptions. However, they rely on few-shot learning with ChatGPT OpenAI (2023) to guide the model’s generation. How to generate reliable data related to molecular knowledge remains a challenge of instruction tuning for existing open source LLMs.
@@ 3.2 Pre-training of PEIT-GEN
Cross-modal causal language modeling. To further strengthen the model’s capability in molecule captioning, we employ the causal language modeling (CLM) to enhance the model performance on text generation. Specifically, we design decoders to generate subsequent property and textual description sequences, under the guidance of SMILES features through cross-attention.
Specifically, given a pair of text and property, the calculation of vanilla self-attentions are as follows:
where () denotes the hidden representations, , , and are the matrix for query, key, and values among the same modality, respectively.
For text decoder and property decoder , we propose cross-modal CLM objectives which further integrates SMILES features for text or property prediction via applying cross-attention:
@@
where we follow SPMM Chang and Ye (2024) to use parameters and for balancing loss terms.
@@ 3.3 Instruction Tuning for PEIT-LLM
Multi-constraint molecule generation task. Molecule generation often requires to be conducted under multiple constraints rather than a single condition. In this work, we propose a new task to assess molecule generation through a variety of descriptors, by comparing the alignment between the generated molecules and specific criteria to evaluate the generative performance of LLMs. By using the large-scale unstructured data generated by PEIT-GEN, we can effectively synthesize sufficient data for evaluation. Specifically, we follow SPMM Chang and Ye (2024) and predict 5 common properties out of the 53 available biochemical properties for diverse SMILES, including BalabanJ, MolLogP, ExactMolWt, QED, and TPSA. Based on the template filling, the predicted multiple property-values can be used to construct data for multi-constraint molecule generation. By using instruction tuning, we guide the LLM to generate molecules while using RDKit Landrum et al. (2013) to verify the actual values of the generated properties. RMSE and R2 are used to compare these values with the constraints to assess the quality of the generated molecules and their alignment with the given conditions. This allows us to systematically evaluate performance of the LLM in multi-constraint molecule generation tasks.
@@
3.4 Comparing PEIT-LLM with Biomolecular Models and LLMs
@@ 4.1 Experimental Setup
Dataset. For pre-training PEIT-GEN, we extract approximately 480k molecular SMILES entries from the ZINC dataset Irwin et al. (2012) and then generate SMILES-text pair data using MolT5 Edwards et al. (2022). Additionally, we calculate 53 biochemical property-value via RDKit, resulting in nearly 480k “text-SMILES-properties” triplets for pre-training. Following MolT5, we use the CHEBI-20 dataset Edwards et al. (2021) to evaluate PEIT-GEN’s performance on molecule captioning and molecular property prediction. We split the CHEBI-20 dataset into training, validation, and test sets with an 8:1:1 ratio, and we verify the property values of each molecule via RDKit. And we use MoleculeNet dataset Wu et al. (2018) to further evaluate the generalization of PEIT-GEN. Details of these datasets are provided in Appendix B.
@@
Baseline Models. We compare our model111Codes can be found in the supplementary materials., PEIT-GEN and PEIT-LLM, against three types of baselines as follows: Baselines on molecule caption such as MolT5 Edwards et al. (2022), BioT5 Pei et al. (2023), MolCA Liu et al. (2023c), Text+Chem-T5 Christofidellis et al. (2023), GIT-Mol Liu et al. (2024). Baselines on moleucular property prediction such as SPMM Chang and Ye (2024), D-MPNN Yang et al. (2019), PretrainGNN Hu et al. (2019), GROVERlarge Rong et al. (2020), ChemRL-GEM Fang et al. (2022). Baselines of LLMs such as LLaMa3 Touvron et al. (2023), Qwen2.5 Yang et al. (2024), Mol-Instructions Fang et al. (2023), InstructMol-GS Cao et al. (2023), BioMedGPT Zhang et al. (2024b). Details of these baselines and evaluation metric are in Appendix C and D, respectively.
@@
Implementation Details. For pre-training PEIT-GEN, the training batch is 16, temperature is 0.07, and the momentum parameter is 0.995 with AdamW optimizer Loshchilov (2017). We pre-train PEIT-GEN with 20 epochs and then fine-tune it on CHEBI-20 training set for 200 epochs, with a learning rate of 5e-4. For supervised fine-tuning PEIT-LLM, we use LLaMa-Factory Zheng et al. (2024) framework and apply LoRA Hu et al. (2022) fine-tuning for 6 epoches with batch size as 3 and learning rate as 5e-5. The parameter size of each component in PEIT-GEN is provided in Table 6 of Appendix E. All experiments are run on NVIDIA 4090 GPUs with 24GB memory.
4.2 Comparing PEIT-GEN with Pre-trained Biomolecular Models
Molecule captioning. Results on CHEBI-20 molecule captioning are shown in Table 2. Our model demonstrates superior performance in generating high-quality and relevant molecular caption. PEIT-GEN achieved the best results in METEOR and ROUGE, and the second-best performance in BLEU-4. Compared to BioT5 which performs the best in BLEU, our approach requires significantly less data. This indicates that using domain-specific models to generate paired data for pre-training is more efficient than single-modality pre-training.
Molecular property prediction. We evaluate the generalization capability of PEIT-GEN on MoleculeNet Wu et al. (2018) benchmarking datasets, and select four widely-used classification tasks for comparison. Results in Table 3 demonstrate that PEIT-GEN achieves superior AUROC on the Clintox dataset compared to specialized models such as MolFormer Ross et al. (2022) and ChemRL-GEM Fang et al. (2022). Additionally, PEIT-GEN shows competitive performance on other subsets while utilizing less pre-training data the further experiment is provided in Table 7 of Appendix F), highlighting the strong generalization ability of PEIT-GEN in molecular property prediction tasks.
4.3 Comparing PEIT-LLM with LLMs
Molecule captioning. As shown in the top of Table 4, the comparison results show that our model outperforms general-purpose LLMs (Qwen-2.5 and LLaMa3.1) as well as Mol-Instructions and BioMedGPT, which were trained using a biochemical information instruction dataset for SFT. PEIT-LLM achieved the second-best performance on the ROUGE metric and demonstrated competitive results compared to InstructMol-GS, which was trained solely on the CHEBI-20 dataset and has a similar parameter scale as our base model. Case study is provided in Table 8 of Appendix H to further illustrate this point.
@@ 4.3 Comparing PEIT-LLM with LLMs
Text-based molecule generation. The results for text-based molecule generation on the CHEBI-20 test set are shown in bottom of Table 4. PEIT-LLM outperforms other baselines in numerical metrics such as BLEU score, Levenshtein Distance, MACCS Fingerprint Similarity, Morgan Fingerprint Similarity, and RDKit Fingerprint Similarity. Meanwhile Mol-Instructions show an advantage in the Validity metric. This indicates that PEIT-LLM, after multi-task instruction fine-tuning, has a strong understanding of the key structural representations of molecules as well as their textual descriptions. Case study is provided in Table 9 of Appendix H to further illustrate this point. This also indirectly validates the effectiveness of the instruction data synthesized by our proposed PEIT-GEN.
@@
Molecular property prediction. For predicting single property, due to the large number of property, we selected two representative ones for prediction. The property ExactMolWt with relatively large numerical values (usually 1001000), and property MolLogP with relatively small numerical values (usually -510) are shown in Table 5. The results show that PEIT-LLM outperforms all other LLMs in predicting specific biochemical properties, demonstrating that PEIT-LLM exhibits strong sensitivity to molecular properties, showing excellent predictive performance for both properties with large numerical values and those with smaller values. This confirms the feasibility of using multi-task SFT to enhance LLMs’ understanding of molecular properties and further validates the reliability of the molecular property instruction dataset. Case study is provided in Table 10 of Appendix H to further illustrate this point.
@@
Multi-constraint molecule generation. Results for our proposed multi-constraint molecule generation task is shown in Table 5. PEIT-LLM surpasses baselines by large margin in both RMSE and R2 metrics. Case study is provided in Table 11 of Appendix H to further illustrate this point. Note that this task requires the model to meet the demands of multiple properties with precise values, placing high demands on the model’s overall understanding capability. General-purpose LLMs, or those not specifically trained for this task, lack the required information storage and fitting abilities. As demonstrated, through our property enhanced instruction tuning, the model gain strong molecular understanding capabilities.
@@ 4.4 Analyses
Ablation study. Figure 2 shows the ablation study of SMILES-text matching loss and cross-modal contrastive loss , which are not considered in SPMM due to the lack of textual description modality ( and are necessary for caption generation via decoders, thus we do not consider them in ablation study). By removing these training objectives, the performance degradation across all metrics, with a more significant decline when both are removed simultaneously. This demonstrates that both and are helpful in cross-modal feature alignment, thereby enhancing the performance of molecule captioning.
Impact of SFT steps. Figure 3 and Figure 4 show the results of PEIT-LLM with different SFT steps. We find that the performance steadily improved at first few epochs, showing that the instruction data is useful for both molecule captioning and multi-constraint molecule generation tasks. The performance gradually saturates around epochs 5-6. This indicates that the LLaMa-7B model achieves optimal performance with 1 million instruction data, and further training might lead to over fitting.
Impact of SFT steps. Figure 3 and Figure 4 show the results of PEIT-LLM with different SFT steps. We find that the performance steadily improved at first few epochs, showing that the instruction data is useful for both molecule captioning and multi-constraint molecule generation tasks. The performance gradually saturates around epochs 5-6. This indicates that the LLaMa-7B model achieves optimal performance with 1 million instruction data, and further training might lead to over fitting.

### v4

@@ Abstract
Xuan Lin1, Long Chen1†, Yile Wang2††thanks: Corresponding author., Xiangxiang Zeng3, Philip S. Yu4 1School of Computer Science, Xiangtan University 2College of Computer Science and Software Engineering, Shenzhen University 3College of Information Science and Engineering, Hunan University 4Department of Computer Science, University of Illinois jack_lin@xtu.edu.cn wangyile@szu.edu.cn psyu@cs.uic.edu
@@ 1 Introduction
Large language models (LLMs) such as GPT-4 OpenAI (2023), PaLM Chowdhery et al. (2023) and LLaMa Touvron et al. (2023); Dubey et al. (2024) have revolutionized the landscape of artificial intelligence and natural language processing, allowing machines to understand and generate human language with remarkable fluency and coherence. Based on encoded world knowledge Petroni et al. (2019) and powerful instruct-following Zhang et al. (2023) capabilities of LLMs, recent work has successfully used LLM for molecular-related tasks, achieving promising results Fang et al. (2023); Zhang et al. (2024a).
Despite their success, as illustrated in Figure 1 (a) and (b), LLMs still struggle with generating molecules under strict property constraints. Even specialized molecular translation models, such as MolT5 Edwards et al. (2022), fail in these tasks (see Figure 1 (c)). While these models effectively capture relationships between molecular text and structure, they lack sufficient understanding of molecular properties, limiting their ability to incorporate property constraints in prompts. This shortcoming restricts their practical utility in applications like drug discovery Zhavoronkov (2018); Elton et al. (2019).
The challenges in addressing such tasks mainly lie in three aspects: (1) Existing studies have revealed the limitations of LLMs in understanding molecular representations Grisoni (2023), making it difficult to handle tasks requiring precise molecular property comprehension; (2) While there are some known SMILES-property pairing data, it often remains limited to predicting a single property and lacks datasets that cover a wide range of properties Wu et al. (2018). Moreover, most of these datasets do not include precisely described textual data, making it challenging to identify accurate tri-modal data pairs Krenn et al. (2020); (3) To our knowledge, there are no suitable datasets or evaluation methods exist for multi-constraint molecule generation using LLMs, which challenges the standardization and assessment of such tasks Jin et al. (2018); Elton et al. (2019).
@@ 2 Related Work
Instruction Tuning. Constructing specialized instruction datasets is an effective way to enable LLMs to better perform molecular-related tasks. For instance, Mol-Instructions Fang et al. (2023) provides a large-scale biomolecular dataset tailored for LLMs, covering a variety of instructions involving small molecules, proteins, and biomolecular texts. ChemDFM Zhao et al. (2025) advances this paradigm by creating a broader dataset spanning molecular structures, reactions, and properties. Its two-stage training—domain pretraining followed by instruction tuning—enhances the model’s chemical understanding and reasoning capabilities. More recently, GeLLM3O Dey et al. (2025) introduced MuMOInstruct, a high-quality dataset focused on multi-property molecule optimization, demonstrating strong generalization across diverse tasks. Despite these advances, generating reliable and scalable molecular instruction data remains a key challenge, particularly for open-source models.
@@ 3.2 Pre-training of PEIT-GEN
Cross-modal Causal Language Modeling. To further strengthen the model’s capability in molecule captioning, we employ the causal language modeling (CLM) to enhance the model performance on text generation. Specifically, we design decoders to generate subsequent property and textual description sequences, under the guidance of SMILES features through cross-attention as show in Figure 3.
@@
where we follow SPMM Chang and Ye (2024) to use parameters and (:=1:5) for balancing loss terms.
@@ 3.3 Instruction Tuning for PEIT-LLM
Multi-constraint Molecule Generation Task. Molecule generation often requires to be conducted under multiple constraints rather than a single condition. In this work, we propose a new task to assess molecule generation through a variety of descriptors, by comparing the alignment between the generated molecules and specific criteria to evaluate the generative performance of LLMs. By using the large-scale unstructured data generated by PEIT-GEN, we can effectively synthesize sufficient data for evaluation. Specifically, we follow SPMM Chang and Ye (2024) iven the vast number of molecular attributes and the complex combinations thereof, analyzing their impact on results across different molecular counts poses a signifi- cant computational challenge due to resource limitations. To address this, we selected representative ADMET Fu et al. (2024) properties—BalabanJ, MolLogP, ExactMolWt, QED, and TPSA—that capture molecular topology, electronic characteristics, and steric effects, while exhibiting low mutual correlation. Based on template filling, the predicted multi-property values are used to construct data for multi-constraint molecule generation. We employ instruction tuning to guide the LLM in generating molecules, and use RDKit Landrum et al. (2013) to calculate the actual property values. RMSE and R2 are then used to compare these values against the constraints, enabling a systematic evaluation of the LLM’s performance in multi-constraint molecule generation tasks.
@@
3.4 Comparing PEIT-LLM with Biomolecular Models and Large Language Models
@@ 4.1 Experimental Setup
Dataset. For pre-training PEIT-GEN, we extract 480k molecular SMILES from the ZINC dataset Irwin et al. (2012) and generate SMILES-text pairs using MolT5 Edwards et al. (2022). We also compute 53 biochemical properties per molecule using RDKit, forming 480k “text-SMILES-properties” triplets for training.
@@
To evaluate PEIT-GEN and PEIT-LLM, we follow MolT5 by using CHEBI-20 Edwards et al. (2021) and MoleculeNet dataset Wu et al. (2018), with the standard splition into training, validation, and test sets with an 8:1:1 ratio. All property values are verified via RDKit. See details in Appendix B.
Baseline Models. We compare our model, PEIT-GEN and PEIT-LLM, against three types of baselines as follows: Baselines on molecule caption such as MolT5 Edwards et al. (2022), BioT5 Pei et al. (2023), MolCA Liu et al. (2023c), Text+Chem-T5 Christofidellis et al. (2023), GIT-Mol Liu et al. (2024). Baselines on moleucular property prediction such as SPMM Chang and Ye (2024), D-MPNN Yang et al. (2019), PretrainGNN Hu et al. (2019), GROVER Rong et al. (2020), ChemRL-GEM Fang et al. (2022). Baselines of LLMs such as LLaMa3 Touvron et al. (2023), Qwen2.5 Yang et al. (2024), GPT3.5-turbo OpenAI (2023) , Mol-Instructions Fang et al. (2023), InstructMol-GS Cao et al. (2023), BioMedGPT Zhang et al. (2024b), ChemDFM Zhao et al. (2025), ChatGLM GLM et al. (2024), Galactica Taylor et al. (2022), Vicuna Chiang et al. (2023). Details of these baselines and evaluation metric are in Appendix C and D, respectively.
@@
Implementation Details. We pre-train PEIT-GEN for 20 epochs using a batch size of 16, temperature , and momentum 0.995 with the AdamW optimizer Loshchilov (2017). Fine-tuning is then performed on the CHEBI-20 training set for 200 epochs with a learning rate of 5e-4. For supervised fine-tuning of PEIT-LLM, we use the LLaMa-Factory Zheng et al. (2024) framework with LoRA Hu et al. (2022) for 6 epochs, a batch size of 3, and learning rate of 5e-5. The total parameter count of the three encoders and two decoders is 533.8M, with an additional 2.2M parameters for other components. Experiments are conducted on NVIDIA 4090 GPUs with 24GB memory.
4.2 Comparing PEIT-GEN with Pre-trained Biomolecular Models
Molecule Captioning. Results on molecule captioning using CHEBI-20 dataset are shown in Table 2. Our model demonstrates superior performance in generating high-quality and relevant molecular caption. PEIT-GEN achieved the best results in METEOR and ROUGE, and the second-best performance in BLEU-4. Compared to BioT5 which performs the best in BLEU, our approach requires significantly less data. This indicates that using domain-specific models to generate paired data for pre-training is more efficient than single-modality pre-training.
Molecular Property Prediction. We evaluate the generalization ability of PEIT-GEN on the MoleculeNet benchmark Wu et al. (2018) using four widely adopted tasks. As shown in Table 3, PEIT-GEN outperforms specialized models such as MolFormer Ross et al. (2022) and ChemRL-GEM Fang et al. (2022) on the Clintox dataset. Despite using less pre-training data, it remains competitive on other subsets. To further demonstrate its predictive strength across 53 molecular properties, we present a relative difference analysis in Appendix E, highlighting PEIT-GEN’s strong generalization in property prediction.
4.3 Comparing PEIT-LLM with LLMs
Molecule Captioning. As shown in the top of Table 4, the comparison results show that our model outperforms general-purpose Qwen-2.5 and LLaMa3.1 as well as Mol-Instructions and BioMedGPT, which were trained using a biochemical information instruction dataset for SFT. PEIT-LLM achieved the second-best performance on the ROUGE metric and demonstrated competitive results compared to InstructMol-GS, which was trained solely on the CHEBI-20 dataset and has a similar parameter scale as our base model.
@@ 4.3 Comparing PEIT-LLM with LLMs
Text-based Molecule Generation. Results on the CHEBI-20 test set are presented at the bottom of Table 4. PEIT-LLM outperforms all baselines on numerical metrics, including BLEU, Levenshtein Distance, and fingerprint similarities based on MACCS, Morgan, and RDKit. Although Mol-Instructions achieves the highest Validity score, the results demonstrate that PEIT-LLM, after multi-task instruction fine-tuning, effectively captures key molecular structures and their corresponding textual representations. A case study in Table 9 of Appendix F further supports these findings and indirectly validates the quality of the instruction data generated by PEIT-GEN.
@@
Molecular Property Prediction. For single-property prediction, due to the large number of available properties, we select two representative examples: ExactMolWt, which typically has large numerical values (100 to 1000), and MolLogP, with smaller values (–5 to 10), as shown in Table 5. The results show that PEIT-LLM consistently outperforms other LLMs in predicting these biochemical properties, demonstrating strong sensitivity and adaptability to molecular property scales. This highlights the effectiveness of multi-task SFT in enhancing LLMs’ understanding of molecular characteristics and further validates the quality and reliability of our molecular property instruction dataset. A case study is provided in Table 10 of Appendix F for further illustration.
@@
Multi-constraint Molecule Generation. Results for our proposed multi-constraint molecule generation task is shown in Table 5. PEIT-LLM surpasses baselines by large margin in both RMSE and R2 metrics. Case study is provided in Table 11 of Appendix F to further illustrate this point. Note that this task requires the model to meet the demands of multiple properties with precise values, placing high demands on the model’s overall understanding capability. General-purpose LLMs, or those not specifically trained for this task, lack the required information storage and fitting abilities. The model gain strong molecular understanding capabilities through property enhanced instruction tuning.
| Model | BLEU↑ | METEOR↑ | ROUGE-2↑ | ROUGE-L↑ |
| Galactica-6.7B Taylor et al. (2022) | 0.008 | 0.065 | 0.015 | 0.063 |
| MolT5-248M Edwards et al. (2022) | 0.001 | 0.033 | 0.001 | 0.034 |
| Vicuna-7B Chiang et al. (2023) | 0.011 | 0.168 | 0.055 | 0.130 |
| Text+Chem T5-223M Christofidellis et al. (2023) | 0.036 | 0.139 | 0.075 | 0.119 |
| ChatGLM-6B GLM et al. (2024) | 0.011 | 0.105 | 0.066 | 0.148 |
| LLaMa3.1-8B Dubey et al. (2024) | 0.014 | 0.184 | 0.066 | 0.148 |
| Qwen2.5-7B Yang et al. (2024) | 0.009 | 0.169 | 0.047 | 0.119 |
| PEIT-LLM-Qwen2.5-7B (Ours) | 0.051 | 0.208 | 0.121 | 0.178 |
| PEIT-LLM-LLaMa3.1-8B (Ours) | 0.053 | 0.215 | 0.125 | 0.184 |
[…]
@@ 4.4 Analyses
Ablation Study. Figure 4 presents an ablation study of the cross-modal matching loss and cross-modal contrastive loss in the PEIT-GEN model for the molecule captioning task ( and are necessary for generation via decoders, thus we do not consider them in ablation study). By removing these training objectives, the performance degradation across all metrics. This demonstrates that both and are helpful in cross-modal feature alignment, thereby enhancing the performance of molecule captioning.
Impact of SFT steps. Figure 5 and Figure 6 illustrate the outcomes of PEIT-LLM across various tasks with different SFT steps. We observe that the performance consistently improves during the initial epochs for all tasks, indicating that the instructional data is beneficial for each, where the performance tends to plateau around epochs 5-6.
Out-of-Distribution Evaluation. To further evaluate the molecular understanding of PEIT-LLM on unseen data, we tested it on the Mol-Instructions Fang et al. (2023) test set, without using the full instructions for pre-training and fine-tuning PEIT-LLMs. As shown in Table 6, PEIT-LLM outperforms all general-purpose LLMs as well as smaller domain-specific models highlighting the strong generalization ability of PEIT-LLM across more diverse molecular instruction tasks.